# ZenSVI: An Open-Source Software for the Integrated Acquisition, Processing and Analysis of Street View Imagery Towards Scalable Urban Science

### v1

@@ Abstract
Street view imagery (SVI) has been instrumental in many studies in the past decade to understand and characterize street features and the built environment. Researchers across a variety of domains, such as transportation, health, architecture, human perception, and infrastructure have employed different methods to analyze SVI. However, these applications and image-processing procedures have not been standardized, and solutions have been implemented in isolation, often making it difficult for others to reproduce existing work and carry out new research. Using SVI for research requires multiple technical steps: accessing APIs for scalable data collection, preprocessing images to standardize formats, implementing computer vision models for feature extraction, and conducting spatial analysis. These technical requirements create barriers for researchers in urban studies, particularly those without extensive programming experience. We develop ZenSVI, a free and open-source Python package that integrates and implements the entire process of SVI analysis, supporting a wide range of use cases. Its end-to-end pipeline includes downloading SVI from multiple platforms (e.g., Mapillary and KartaView) efficiently, analyzing metadata of SVI, applying computer vision models to extract target features, transforming SVI into different projections (e.g., fish-eye and perspective) and different formats (e.g., depth map and point cloud), visualizing analyses with maps and plots, and exporting outputs to other software tools. We demonstrate its use in Singapore through a case study of data quality assessment and clustering analysis in a streamlined manner. Our software improves the transparency, reproducibility, and scalability of research relying on SVI and supports researchers in conducting urban analyses efficiently. Its modular design facilitates extensions and unlocking new use cases. This package is openly available at https://github.com/koito19960406/ZenSVI, and it is supported by documentation including tutorials.
@@ 1 Introduction
Street view imagery (SVI) is a popular geospatial dataset in various fields pertaining to urban science, such as transportation, health, morphology, and urban visual perception, thanks to its scalability, coverage, and spatiotemporal granularity (Biljecki and Ito, 2021; Rzotkiewicz et al., 2018; Kang et al., 2020; Ito et al., 2024b; Wang et al., 2024b). Machine learning techniques to process image data have enabled automated assessments of SVI at large scales, and such scalability has been leveraged with the growing coverage of SVI data platforms. Also importantly, SVI’s ability to capture street-level information and its spatiotemporal granularity has supported many micro-scale assessments of streets and neighborhoods over time that were difficult or impossible with conventional spatial data, such as remote sensing imagery and field surveys (Fan et al., 2023; Kang et al., 2020; Wang et al., 2024c; Liang et al., 2023).
The strengths of SVI over other forms of geospatial data have enabled and supported many applications and use cases for urban studies and data-driven urban planning, and entrenched its importance and role for the foreseeable future. One such example is the development of indexes that quantify diverse urban phenomena. An early instance is the green view index advocated by researchers such as Li et al. (2015) and Yang et al. (2009), which is based on the proportion of greenery in a street view image detected with image segmentation, a computer vision technique. Similarly, other indexes were developed, such as the sky view factor and building view factor (Liang et al., 2017; Wang et al., 2022b; Biljecki et al., 2023). Recently, more comprehensive indexes — the walkability and bikeability index — have been proposed by combining various indicators extracted from SVI through advanced computer vision techniques coupled with geospatial operations (Koo et al., 2022; Ito and Biljecki, 2021). These indicators and indexes derived from SVI have been utilized as explanatory variables to analyze urban phenomena such as land use patterns, housing prices (Law et al., 2019; Zhao et al., 2023a), transportation flow (Jiao and Wang, 2023), socioeconomic status (Fan et al., 2023), and many more (Ning et al., 2022; Cao et al., 2018; Hu et al., 2024; Fujiwara et al., 2024).
In contrast to using pre-computed indicators, direct analysis of SVI through computer vision techniques has enabled various applications such as road condition assessment (Ai et al., 2018; Shu et al., 2021; de Mesquita et al., 2022; Arya et al., 2021); building type and age classification (Kang et al., 2018; Li et al., 2018b; Sun et al., 2022); and accessibility assessments (Weld et al., 2019). Moreover, SVI provides significant insights into assessing people’s subjective perceptions of urban space, such as perceptions of safety (Kang et al., 2023), sound intensity (Zhao et al., 2023b), architectural design (Liang et al., 2024) and attractiveness (Dubey et al., 2016; Zhang et al., 2018). By integrating multi-dimensional features of the built environment captured by SVI, urban planners and researchers can effectively sense the dynamics of urban areas contributing to improving the quality of life for citizens, thereby leading to designing and managing sustainable and resilient cities.
@@
As much as SVI has been a popular and powerful instrument in sensing the built environment, there remain technical difficulties in conducting studies relying on it, especially for researchers without advanced analytical and technical skills. In fact, such difficulties are even on the rise, as large-scale big data analyses with deep learning models are becoming the norm (Ito et al., 2024b). As a consequence, domain experts who are less technical find it challenging to conduct such studies, which hampers equity and participation in research (Ito and Biljecki, 2021; Bromm et al., 2020; Zhu et al., 2025), and may also impede the development of new use cases and the representation of certain geographical areas. Further, most publications are not accompanied with openly released code to allow reproducibility, compounding the issue (Biljecki and Ito, 2021).
@@
OSMnx by Boeing (2017) is a quintessential example — it is a Python package that has facilitated many applications that require network analysis or OpenStreetMap data by integrating them as a one-stop software. Trackintel by Martin et al. (2023) is also a Python package and provides many functions of GPS-trajectory data analysis with standardized models, thereby enhancing the transparency, transferability, and reproducibility of research. Another notable Python package is Urbanity developed by Yap et al. (2023), which allows users to download ready-to-use feature-rich network data. There have also been domain-specific tools, e.g., an R package by Mahajan (2024) computes the green index based on multiple sources of data, such as OpenStreetMap and SVI.
As Yap et al. (2022) highlighted in their comprehensive systematic review, there is still a need for integrative open-source tools that can comparatively evaluate urban environments at a large scale and do so with a relative ease of use and accessibility. This need is particularly acute in SVI research, where even though researchers occasionally share their code, these contributions often address specific niche problems, leading to a fragmented landscape of tools that are difficult to adapt or integrate into broader workflows. Despite the increasing number of applications, trends, and challenges mentioned above, there has been only limited effort to standardize, facilitate, or integrate the workflow of SVI analysis, i.e., an integrated open-source solution to support urban studies relying on SVI.
Therefore, to bridge such a clear gap and facilitate reproducible SVI research, we introduce ZenSVI — a framework and implementation that integrates downloading, processing, transforming, and visualizing SVI, which will benefit and encourage studies by improving the transparency, methodology, and ease of conducting research that requires SVI and associated elements such as computer vision models. This study is the first to create such an integrated software not only in SVI research but also in urban imagery. ZenSVI, our open-source Python package based on open data and open models, will further accelerate innovative urban science research by mitigating barriers for less experienced researchers and making their research reproducible. On a broader scope, it presents a novel contribution to the open science current of urban analytics, and its modular framework may contribute to its extensions or integrations in other workflows.
@@ 2.1 Open-source software for urban science
raster visionh Satellite imagery $\bullet$
Open-source software tools are increasingly used in urban studies, enabling researchers to analyze complex urban phenomena efficiently and collaboratively (Yap et al., 2022; Pönkänen et al., 2025). These tools have contributed to advancing urban science by providing freely and easily accessible platforms for acquiring, processing, analyzing, and visualizing urban data. For example, OSMnx has facilitated the study of urban networks by streamlining the extraction, construction, and analysis of street networks from OpenStreetMap. Such tools allow researchers to tackle large-scale spatial analyses with minimal technical barriers.
The strength of these open-source tools lies not in introducing new scientific methods, but rather in their ability to democratize existing methodologies and foster collaborative improvement through community engagement. Their success stems from implementing established scientific approaches while remaining open to enhancement by a community of like-minded developers and researchers. These tools excel at facilitating essential tasks such as downloading large datasets and visualizing complex spatial information, often serving as comprehensive one-stop solutions for specific analytical needs. Additionally, some have integrated machine learning (ML) approaches for tasks such as feature extraction without needing to build models from scratch (Table 1). By embedding cutting-edge algorithms and workflows in user-friendly interfaces, these tools empower researchers across disciplines, including those without extensive programming expertise, to conduct innovative studies. This inclusivity expands the reach and impact of urban science.
@@ 2.2 Open-source software for street view imagery analysis
SVI has a wide range of applications in urban studies (Biljecki and Ito, 2021), including urban planning and management (Wang et al., 2023; Li et al., 2022c; Xia et al., 2021a; Lu, 2019; Maniat et al., 2021; Wang et al., 2022b), environmental monitoring and assessment (Cao et al., 2018; He and Li, 2021; Xia et al., 2021a; Li et al., 2022d; Yin and Wang, 2016; Chen and Biljecki, 2023; Li et al., 2022b; Liang et al., 2023; Lu et al., 2023; Sanusi et al., 2016; Li et al., 2018a), transportation research (Campbell et al., 2019; Cai et al., 2022), and socio-economic studies (Alhasoun and González, 2019; Zhang and Dong, 2018). Although these applications solve specific problems, they share many of the same approaches and data sources, but no effort has been made to integrate these applications into one platform so that other researchers can reproduce their work easily.
@@
SVI data have diverse sources and acquisition methods. The most widely used SVI comes from commercial GIS service platforms such as Google [^1] https://www.google.com/maps, Amap [^2] https://www.amap.com/ and Baidu [^3] https://map.baidu.com/, which capture high-resolution, geotagged panoramic images via vehicles (Mahabir et al., 2020; Chen et al., 2019). These platforms offer extensive coverage suitable for urban research but are subject to terms of use and privacy policies (Hou and Biljecki, 2022; Helbich et al., 2024). Open projects like KartaView and Mapillary consolidate and provide user-uploaded SVI, offering free, scalable alternatives, though often with lower consistency and quality (Hara et al., 2013; Zheng and Amemiya, 2023; Juhász and Hochmair, 2016; Hou et al., 2024). Additionally, some city governments and public institutions collect and release SVI, and social media platforms (e.g., Flickr and Instagram) are also significant sources of imagery (Bahrehdar et al., 2020; Chen et al., 2020b; Callau et al., 2019).
@@
Processing SVI typically involves complex computer vision techniques to extract and analyze valuable information, broadly categorized into traditional methods and deep learning-based approaches. The former includes edge detection, texture analysis, color feature extraction, and shape feature extraction (Bansal et al., 2011; de Souza-Daw et al., 2015). With advancements in deep learning, methods such as object detection, semantic segmentation, and image classification have become predominant. Object detection identifies and locates objects (e.g., vehicles, pedestrians, traffic signs) using models such as YOLO, Faster R-CNN, and R-CNN (Krylov et al., 2018; Campbell et al., 2019; Kang et al., 2020; Ren et al., 2023; Zhang et al., 2020). Semantic segmentation assigns pixel-level labels (e.g., roads, buildings, vegetation) via models such as U-Net, SegNet, and DeepLab (Xia et al., 2021b; Aikoh et al., 2023; Wang et al., 2022a; Gou et al., 2022; Zhao et al., 2023b). Image classification categorizes entire images (e.g., commercial, residential, industrial) using models such as ResNet and VGG (Law et al., 2020; Fang et al., 2022; Huang et al., 2024).
@@ 3 ZenSVI framework and implementation
ZenSVI is conceived as a one-stop open-source solution for SVI analytics, built upon robust scientific and software engineering practices to ensure maintainability, extensibility, and code quality. The package follows a modular architecture with a clear separation of concerns, extensive unit testing, continuous integration through GitHub Actions, and comprehensive API documentation. Further, as an open-source project, ZenSVI welcomes contributions from the community, enabling researchers and developers to extend its functionality, improve existing features, or address specific use cases. This collaborative development approach ensures that the package can evolve alongside the changing needs of the urban research community while maintaining high standards of code quality and documentation.
The framework is designed with compatibility in mind, leveraging widely used Python packages in the geospatial ecosystem. Core dependencies include GeoPandas [^8] https://github.com/geopandas/geopandas for spatial operations (Jordahl et al., 2020), OSMnx for street network analysis, and PyTorch for deep learning implementations. This foundation ensures seamless integration with existing geospatial workflows while maintaining forward compatibility through semantic versioning. The framework’s reliance on these well-maintained, community-supported packages helps ensure long-term continuity and stability. Regular updates are guaranteed through continuous integration and automated testing workflows, while the modular architecture allows for easy integration of new features and improvements without disrupting existing functionality.
For computer vision capabilities, we integrate established models from other research efforts, maintaining their original architectures and parameters to ensure reproducibility and validated performance. Each implemented model is thoroughly documented with citations to its source publication, allowing researchers to reference the original methodologies and performance metrics.
@@
To ensure sustainable development, we maintain detailed contribution guidelines, issue templates, and a roadmap for future development on GitHub. The framework is designed for both direct usage and extension through clear interfaces and plugin architecture. We actively encourage community contributions while maintaining code quality through automated checks and peer review processes.
@@ 3.1 Download
Built on powerful geospatial libraries such as GeoPandas and OSMnx (Boeing, 2017), this sub-package implements efficient spatial processing capabilities specifically designed for acquiring SVI through wrapper functions for two major platforms: Mapillary and KartaView, which have global relevance. In addition, as a demonstration of further and localized capabilities, and in recognition of accelerating city and national government efforts in acquiring SVI, we support Amsterdam’s Government-hosted API.
The package standardizes input types and output formats while enhancing the official APIs’ functionalities. Users can provide location data for SVI acquisition in various formats: coordinate pairs, CSV files with longitude and latitude columns, shapefiles, GeoJSON files, or place names as string values. By leveraging GeoPandas for spatial operations and OSMnx for street network analysis, the package efficiently processes these inputs to generate and filter sampling points for SVI downloads, handling queries at various scales, from neighborhood to global level.
To ensure robust SVI acquisition, the download function implements parallel processing of points and incorporates checkpoints for handling large-scale downloads. This parallel computing approach significantly improves the speed and reliability of acquiring data, particularly when downloading millions of images across extensive geographic areas. These enhancements, built upon established spatial libraries, streamline the user experience by eliminating the need for custom API integration code while maintaining reproducibility in the SVI acquisition process.

### v2

@@ Abstract
Street view imagery (SVI) has been instrumental in many studies in the past decade to understand and characterize street features and the built environment. Researchers across a variety of domains, such as transportation, health, architecture, human perception, and infrastructure have employed different methods to analyze SVI. However, these applications and image-processing procedures have not been standardized, and solutions have been implemented in isolation, often making it difficult for others to reproduce existing work and carry out new research. Using SVI for research requires multiple technical steps: accessing APIs for scalable data collection, preprocessing images to standardize formats, implementing computer vision models for feature extraction, and conducting spatial analysis. These technical requirements create barriers for researchers in urban studies, particularly those without extensive programming experience. We developed ZenSVI, a free and open-source Python package that integrates and implements the entire process of SVI analysis, supporting a wide range of use cases. Its end-to-end pipeline includes downloading SVI from multiple platforms (e.g., Mapillary and KartaView) efficiently, analyzing metadata of SVI, applying computer vision models to extract target features, transforming SVI into different projections (e.g., fish-eye and perspective) and different formats (e.g., depth map and point cloud), visualizing analyses with maps and plots, and exporting outputs to other software tools. We demonstrated its use in Singapore through a case study of data quality assessment and clustering analysis in a streamlined manner. Our software improves the transparency, reproducibility, and scalability of research relying on SVI and supports researchers in conducting urban analyses efficiently. Its modular design facilitates extensions of the package for new use cases. This package is openly available at https://github.com/koito19960406/ZenSVI, and it is supported by documentation including tutorials (https://zensvi.readthedocs.io/en/latest/examples/index.html).
@@ 1 Introduction
Street view imagery (SVI) is a popular geospatial dataset in various fields pertaining to urban science, such as transportation, health, morphology, and urban visual perception, thanks to its scalability, coverage, and spatiotemporal granularity (Biljecki and Ito, 2021; Rzotkiewicz et al., 2018; Kang et al., 2020; Ito et al., 2024b; Wang et al., 2024b). Machine learning techniques to process image data have enabled automated assessments of SVI on large scales, and such scalability has been leveraged with the growing coverage of SVI data platforms. Also importantly, SVI’s ability to capture street-level information and its spatiotemporal granularity has supported many micro-scale assessments of streets and neighborhoods over time that were difficult or impossible with conventional spatial data, such as remote sensing imagery and field surveys (Fan et al., 2023; Kang et al., 2020; Wang et al., 2024c; Liang et al., 2023).
The strengths of SVI have enabled numerous applications in urban studies and data-driven planning. Researchers have developed various indices to quantify urban phenomena, from the green view index (Li et al., 2015; Yang et al., 2009) to sky view and building view factors (Liang et al., 2017; Wang et al., 2022b; Biljecki et al., 2023), and comprehensive walkability and bikeability indices (Koo et al., 2022; Ito and Biljecki, 2021). These SVI-derived indicators have been used to analyze diverse urban phenomena including land use patterns, housing prices (Law et al., 2019; Zhao et al., 2023a), transportation flows (Jiao and Wang, 2023), and socioeconomic status (Fan et al., 2023; Ning et al., 2022; Cao et al., 2018; Hu et al., 2024; Fujiwara et al., 2024). SVI has also been utilized to predict or estimate objective and subjective characteristics of streets, such as road condition assessment (Ai et al., 2018; Shu et al., 2021; de Mesquita et al., 2022; Arya et al., 2021), building classification (Kang et al., 2018; Li et al., 2018b; Sun et al., 2022), accessibility evaluation (Weld et al., 2019), and assessment of subjective perceptions including safety (Kang et al., 2023), sound intensity (Zhao et al., 2023b), architectural design (Liang et al., 2024), and attractiveness (Dubey et al., 2016; Zhang et al., 2018).
@@
As much as SVI has been a popular and powerful instrument in sensing the built environment, there remain technical difficulties in conducting studies relying on it, especially for researchers without advanced analytical and technical skills. In fact, such difficulties have even been on the rise, as large-scale big data analyses with deep learning models are becoming the norm (Ito et al., 2024b). As a consequence, domain experts who are less technical find it challenging to conduct such studies, which hampers equity and participation in research (Ito and Biljecki, 2021; Bromm et al., 2020; Zhu et al., 2025), and may also impede the development of new use cases and the representation of certain geographical areas. Further, most publications are not accompanied by openly released code to allow reproducibility, compounding the issue (Biljecki and Ito, 2021).
@@
OSMnx by Boeing (2017) is a quintessential example — it is a Python package that has facilitated many applications that require network analysis or OpenStreetMap data by integrating them as a one-stop software. Trackintel by Martin et al. (2023a) is also a Python package and provides many functions of GPS-trajectory data analysis with standardized models, thereby enhancing the transparency, transferability, and reproducibility of research. Another notable Python package is Urbanity developed by Yap et al. (2023), which allows users to download ready-to-use feature-rich network data. There have also been domain-specific tools, e.g., an R package by Mahajan (2024) computes the green index based on multiple sources of data, such as OpenStreetMap and SVI.
As Yap et al. (2022) highlighted in their comprehensive systematic review, there is still a need for integrative open-source tools that can comparatively evaluate urban environments on a large scale and do so with relative ease of use and accessibility. This need is particularly acute in SVI research, where even though researchers occasionally share their code, these contributions often address specific niche problems, leading to a fragmented landscape of tools that are difficult to adapt or integrate into broader workflows. Despite the increasing number of applications, trends, and challenges mentioned above, researchers have only made a limited effort to standardize, facilitate, or integrate the workflow of SVI analysis, i.e., an integrated open-source solution to support urban studies relying on SVI.
Therefore, to bridge such a clear gap and facilitate reproducible SVI research, we introduce ZenSVI — a framework and implementation that integrates downloading, processing, transforming, and visualizing SVI, which will benefit and encourage studies by improving the transparency, methodology, and ease of conducting research that requires SVI and associated elements such as computer vision models. This study is the first to create such an integrated software not only in SVI research but also in urban imagery. ZenSVI simplifies many layers of SVI analysis, such as data acquisition and processing, and will therefore accelerate innovative urban science research by mitigating barriers for less experienced researchers and making their research reproducible. In a broader scope, its framework based on open data and open models presents a novel contribution to the open science current of urban analytics, and its modular framework may contribute to its extensions or integrations in other workflows.
@@ 2.1 Open-source software for urban science
Trackintel (Martin et al., 2023b) h Mobility $\bullet$ $\bullet$
GWmodelS (Lu et al., 2023a) i Spatial analysis $\bullet$ $\bullet$
raster visionj Satellite imagery $\bullet$
Open-source software tools are increasingly used in urban studies, enabling researchers to analyze complex urban phenomena efficiently and collaboratively (Yap et al., 2022; Pönkänen et al., 2025). These tools have contributed to the advancement of urban science by providing freely and easily accessible platforms for acquiring, processing, analyzing, and visualizing urban data. For example, OSMnx has facilitated the study of urban networks by streamlining the extraction, construction, and analysis of street networks from OpenStreetMap. Such tools allow researchers to tackle large-scale spatial analyses with minimal technical barriers.
The strength of these open-source tools lies not in introducing new scientific methods, but rather in their ability to democratize existing methodologies and foster collaborative improvement through community engagement. Their success stems from implementing established scientific approaches while remaining open to enhancement by a community of like-minded developers and researchers. These tools excel at facilitating essential tasks, such as downloading large datasets and visualizing complex spatial information, often serving as comprehensive one-stop solutions for specific analytical needs. Additionally, some have integrated machine learning (ML) approaches for tasks such as feature extraction without the need to build models from scratch (Table 1). By embedding cutting-edge algorithms and workflows in user-friendly interfaces, these tools empower researchers across disciplines, including those without extensive programming experience, to conduct innovative studies. This inclusivity expands the reach and impact of urban science.
@@ 2.2 Open-source software for street view imagery analysis
SVI has a wide range of applications in urban studies (Biljecki and Ito, 2021), including urban planning and management (Wang et al., 2023; Li et al., 2022c; Xia et al., 2021a; Lu, 2019; Maniat et al., 2021; Wang et al., 2022b), environmental monitoring and evaluation (Cao et al., 2018; He and Li, 2021; Xia et al., 2021a; Li et al., 2022d; Yin and Wang, 2016; Chen and Biljecki, 2023; Li et al., 2022b; Liang et al., 2023; Lu et al., 2023b; Sanusi et al., 2016; Li et al., 2018a), transportation research (Campbell et al., 2019; Cai et al., 2022), and socioeconomic studies (Alhasoun and González, 2019; Zhang and Dong, 2018). Although these applications solve specific problems, they share many of the same approaches and data sources, but no effort has been made to integrate these applications into one platform so that other researchers can reproduce their work easily.
@@
SVI data have diverse sources and acquisition methods. The most widely used SVI comes from commercial GIS service platforms such as Google [^1] https://www.google.com/maps, Amap [^2] https://www.amap.com/ and Baidu [^3] https://map.baidu.com/, which capture high-resolution geotagged panoramic images via vehicles (Mahabir et al., 2020; Chen et al., 2019). These platforms offer a wide coverage suitable for urban research but are subject to terms of use and privacy policies (Hou and Biljecki, 2022; Helbich et al., 2024). Open projects like KartaView and Mapillary consolidate and provide user-uploaded SVI, offering free, scalable alternatives, although often with lower consistency and quality (Hara et al., 2013; Zheng and Amemiya, 2023; Juhász and Hochmair, 2016; Hou et al., 2024). Additionally, some city governments and public institutions collect and release SVI, and social media platforms (e.g., Flickr and Instagram) are also significant sources of imagery (Bahrehdar et al., 2020; Chen et al., 2020b; Callau et al., 2019).
@@
Processing SVI typically involves complex computer vision techniques to extract and analyze valuable information, broadly categorized into traditional methods and deep learning-based approaches. The former includes edge detection, texture analysis, color feature extraction, and shape feature extraction (Bansal et al., 2011; de Souza-Daw et al., 2015). With advancements in deep learning, object detection, semantic segmentation, and image classification have become predominant. Object detection identifies and locates objects (e.g., vehicles, pedestrians, traffic signs) using models such as YOLO, Faster R-CNN, and R-CNN (Krylov et al., 2018; Campbell et al., 2019; Kang et al., 2020; Ren et al., 2023; Zhang et al., 2020). Semantic segmentation assigns pixel-level labels (e.g., roads, buildings, vegetation) via models such as U-Net, SegNet, and DeepLab (Xia et al., 2021b; Aikoh et al., 2023; Wang et al., 2022a; Gou et al., 2022; Zhao et al., 2023b). Image classification categorizes entire images (e.g., commercial, residential, industrial) using models such as ResNet and VGG (Law et al., 2020; Fang et al., 2022; Huang et al., 2024).
@@ 3 ZenSVI framework and implementation
ZenSVI is conceived as a one-stop open-source solution for SVI analytics and is built with a modular architecture to ensure maintainability, extensibility, and code quality. The package is developed with comprehensive unit tests, automated continuous integration/continuous deployment (CI/CD) pipelines through GitHub Actions and detailed API documentation with usage examples. Further, as an open-source project, ZenSVI welcomes contributions from the community, enabling researchers and developers to extend its functionality, improve existing features, or address specific use cases. This continuous and collaborative development approach ensures that the package can evolve alongside the changing needs of the urban research community while maintaining high standards of code quality and documentation.
The framework is designed with compatibility in mind, leveraging widely used Python packages in the geospatial ecosystem. Core dependencies include GeoPandas [^8] https://github.com/geopandas/geopandas for spatial operations (Jordahl et al., 2020), OSMnx for street network analysis, and PyTorch for deep learning implementations. This foundation ensures seamless integration with existing geospatial workflows while maintaining forward compatibility through semantic versioning. The framework’s reliance on these well-maintained, community-supported packages helps ensure long-term continuity and stability.
For computer vision capabilities, we integrated established models from other research efforts, maintaining their original architectures and parameters to ensure reproducibility and validated performance. Also, each implemented model is documented with citations to its source publication.
@@ 3.1 Download
Built on geospatial libraries such as GeoPandas and OSMnx (Boeing, 2017), this sub-package implements wrapper functions with parallel processing to acquire SVI from two major platforms: Mapillary and KartaView, which have global coverage. In addition, the City of Amsterdam’s SVI API is supported, which showcases ZenSVI’s capability to accommodate various data sources, not limited to commercial or crowdsourced platforms.
The package standardizes input types and output formats while enhancing the official APIs’ functionalities. Users can provide location data for SVI acquisition in various formats: coordinate pairs, CSV files with longitude and latitude columns, shapefiles, GeoJSON files, or place names as string values. By leveraging GeoPandas for spatial operations and OSMnx for street network analysis, the package efficiently and robustly processes these inputs to generate and filter sampling points for SVI downloads with parallel processing and checkpoints. With these additional features, ZenSVI eliminates the need for custom API integration code by researchers.